# Understanding Deep Representation Learning via Layerwise Feature Compression and Discrimination

### v3

@@ B.5 Proof of 1
Proof.
Since GD converges to a global optimal solution, we have $\bm{W}_{L:1}\bm{X}=\bm{Y}$. This, together with the fact that $\bm{X}$ is square and orthogonal, yields (10). According to [31, 3], the iterates of gradient flow for solving Problem (3) satisfy (19). This, together with the initialization (18), yields for all $t\geq 0$,
@@
This implies that (11) holds with $\delta=\xi^{2}\sqrt{d-K}$. Moreover, using [124, Theorem 1], we obtain that (12) holds with $\varepsilon=\xi$ and $\rho=0$. Then, the proof is completed. ∎
@@
Figure 14: Approximate low-rankness hold universally across different training and network configurations. The experimental setup is deferred to Section 5.2.1. For each figure, the vertex of each polygon denotes the singular value of the weights for each $d$ and $\xi$. As we observe, for each fixed $\xi$, the polygon is a hexagon, implying that the singular values are approximately equal and independent of the network width $d$. Second, the singular values grow when the initialization scale $\xi$ increases and also confirms 1 that $\varepsilon=\xi$. Moreover, we conclude that $\sigma_{\min}\approx\sigma_{\max}$ when comparing (a) and (b). Therefore, we can see that both the minimum and maximum singular values remain unchanged, demonstrating that the GD operates only within a 2K-dimensional invariant subspace.
Appendix C Extra Experiments
@@
C.1 Assumption Discussion
Here, we provide more experimental results for validating the low-rankness property in 2 of trained weights in DLN. To validate the low-rankness assumption, we train 4-layer DLNs using the orthogonal initialization with width $d\in\left\{50,100,200,300,400,500\right\}$ and initialization scale $\xi\in\left\{0.1,0.2,0.3,0.4,0.5\right\}$. The training setting is the same with Section 5.2. For the weights $\left\{\bm{W}_{l}\right\}_{l=1}^{L-1}$ of each network that have been trained until convergence, we plot
minimum singular $\text{minimum singular value}\quad\sigma_{\min}\;$ $:=\;\min_{K+1\leq i\leq d-K,\;l\in[L-1]}\sigma_{i}(\bm{W}_{l}),$
maximum singular $\text{maximum singular value}\quad\sigma_{\max};$ $:=\;\max_{K+1\leq i\leq d-K,\;l\in[L-1]}\sigma_{i}(\bm{W}_{l}).$
The results are shown in Figure 14. Here, for each $l\in[L-1]$, we assume that $\sigma_{1}(\bm{W}_{l})\geq\sigma_{2}(\bm{W}_{l})\geq\cdots\geq\sigma_{d}(\bm{W}_{l})$. For each figure, the vertex of each polygon denotes the singular value of the weights for each $d$ and $\xi$. As we observe, for each fixed $\xi$, the polygon is a hexagon, implying that the singular values are approximately equal and independent of the network width $d$. Second, the singular values grow when the initialization scale $\xi$ increases and also confirms 1 that $\varepsilon=\xi$. Moreover, we conclude that $\sigma_{\min}\approx\sigma_{\max}$ when comparing Figure 14(a) and Figure 14(b). Therefore, we can see that both the minimum and maximum singular values remain unchanged, demonstrating that the GD operates only within a 2K-dimensional invariant subspace.
C.2 Exploration on Other Datasets
@@
To further illustrate the analogous roles of deep linear and nonlinear layers across different data modalities, we reproduce the experiments from Figure 1 using the SST-5 text dataset [102], keeping all experimental settings unchanged except for the dataset. As shown in Figure 15, we observe a similar trend: early nonlinear layers enhance linear separability, while deeper layers—regardless of their nonlinearity—progressively compress the features. We also use this experimental setup to verify 1, with results presented in Figure 16. As observed, within-class compression and between-class discrimination respectively exhibit progressively collapsing and increasing trends, further confirming that our theoretical insights generalize beyond the image domain.
@@
Due to resource constraints, we use the PyTorch pre-trained VGG11 network [108] on ImageNet [28] to extract features and evaluate within-class compression and between-class discrimination. The results are presented in Figure 17. We observe that the compression metric continues to exhibit an approximately geometric decay, and the discrimination metric shows increasing pattern, thereby validating our theory on the large-scale benchmark dataset.

### v4

@@ A.5 Proof of Proposition 5
Proof Since GD converges to a global optimal solution, we have $\bm{W}_{L:1}\bm{X}=\bm{Y}$. This, together with the fact that $\bm{X}$ is square and orthogonal, yields (10). According to du2018algorithmic; arora2018optimization, the iterates of gradient flow for solving Problem (3) satisfy (19). This, together with the initialization (18), yields for all $t\geq 0$,
@@
This implies that (11) holds with $\delta=\xi^{2}\sqrt{d-K}$. Moreover, using (yaras2023law, Theorem 1), we obtain that (12) holds with $\varepsilon=\xi$ and $\rho=0$. Then, the proof is completed.
@@
Figure 14: Approximate low-rankness hold universally across different training and network configurations. The experimental setup is deferred to Section 5.2.1. For each figure, the vertex of each polygon denotes the singular value of the weights for each $d$ and $\xi$. As we observe, for each fixed $\xi$, the polygon is a hexagon, implying that the singular values are approximately equal and independent of the network width $d$. Second, the singular values grow when the initialization scale $\xi$ increases and also confirms Proposition 5 that $\varepsilon=\xi$. Moreover, we conclude that $\sigma_{\min}\approx\sigma_{\max}$ when comparing (a) and (b). Therefore, we can see that both the minimum and maximum singular values remain unchanged, demonstrating that the GD operates only within a 2K-dimensional invariant subspace.
Appendix B Extra Experiments
@@
B.1 Assumption Discussion
Here, we provide more experimental results for validating the low-rankness property in 3 of trained weights in DLN. To validate the low-rankness assumption, we train 4-layer DLNs using the orthogonal initialization with width $d\in\{50,100,200,300,400,500\}$ and initialization scale $\xi\in\{0.1,0.2,0.3,0.4,0.5\}$. The training setting is the same with Section 5.2. For the weights $\{\bm{W}_{l}\}_{l=1}^{L-1}$ of each network that have been trained until convergence, we plot
$\text{minimum singular value}\quad\sigma_{\min}\;$ $:=\;\min_{K+1\leq i\leq d-K,\;l\in[L-1]}\sigma_{i}(\bm{W}_{l}),$
$\text{maximum singular value}\quad\sigma_{\max};$ $:=\;\max_{K+1\leq i\leq d-K,\;l\in[L-1]}\sigma_{i}(\bm{W}_{l}).$
The results are shown in Figure 14. Here, for each $l\in[L-1]$, we assume that $\sigma_{1}(\bm{W}_{l})\geq\sigma_{2}(\bm{W}_{l})\geq\cdots\geq\sigma_{d}(\bm{W}_{l})$. For each figure, the vertex of each polygon denotes the singular value of the weights for each $d$ and $\xi$. As we observe, for each fixed $\xi$, the polygon is a hexagon, implying that the singular values are approximately equal and independent of the network width $d$. Second, the singular values grow when the initialization scale $\xi$ increases and also confirms Proposition 5 that $\varepsilon=\xi$. Moreover, we conclude that $\sigma_{\min}\approx\sigma_{\max}$ when comparing Figure 14(a) and Figure 14(b). Therefore, we can see that both the minimum and maximum singular values remain unchanged, demonstrating that the GD operates only within a 2K-dimensional invariant subspace.
B.2 Exploration on Other Datasets
@@
To further illustrate the analogous roles of deep linear and nonlinear layers across different data modalities, we reproduce the experiments from Figure 1 using the SST-5 text dataset (socher2013recursive), keeping all experimental settings unchanged except for the dataset. As shown in Figure 15, we observe a similar trend: early nonlinear layers enhance linear separability, while deeper layers—regardless of their nonlinearity—progressively compress the features. We also use this experimental setup to verify Theorem 4, with results presented in Figure 16. As observed, within-class compression and between-class discrimination respectively exhibit progressively collapsing and increasing trends, further confirming that our theoretical insights generalize beyond the image domain.
@@
Due to resource constraints, we use the PyTorch pre-trained VGG11 network (simonyan2014very) on ImageNet (deng2009imagenet) to extract features and evaluate within-class compression and between-class discrimination. The results are presented in Figure 17. We observe that the compression metric continues to exhibit an approximately geometric decay, and the discrimination metric shows increasing pattern, thereby validating our theory on the large-scale benchmark dataset.
@@
Appendix C Auxiliary Results
Lemma 14
Proof Let $\bm{A}=\bm{U}\bm{\Sigma}\bm{V}^{T}$ be a singular value decomposition of $\bm{A}$, where
$\bm{U}\in\mathcal{O}^{m}$, and $\bm{V}\in\mathcal{O}^{n}$. Then we compute
which, together with the AM-QM inequality, directly implies (108).
Lemma 15
Lemma 16
$\bm{Q}\bm{Q}^{T}=\bm{U}\bm{U}^{T}.$ (110)
Proof It follows from $\bm{U}\bm{\Lambda}\bm{U}^{T}=\bm{Q}\bm{A}\bm{Q}^{T}$, $\bm{Q}\in\mathcal{O}^{d\times K}$, and $\bm{U}\in\mathcal{O}^{d\times K}$ that
$\bm{Q}^{T}\bm{U}\bm{\Lambda}=\bm{A}\bm{Q}^{T}\bm{U}.$ (111)
Now, left-multiplying both sides by $\bm{Q}$ yields
Lemma 17
For every real number $1\leq\theta\leq 1$ and $x\geq-1$, it holds that
$(1+x)^{\theta}\leq 1+\theta x.$